OpenStudy (anonymous):

for (x+6) (4x+1) use the discriminant to describe the type of solutions

OpenStudy (anonymous):

you don't have to because the solution is x = -6 or x = -1/4

OpenStudy (mertsj):

If the directions say to use the discriminant then that is what the student should do.

OpenStudy (anonymous):

How can I find the discriminant for these factors?

OpenStudy (anonymous):

you need to foil

OpenStudy (mathmale):

My perspective is that if you have two real roots to your quadratic equation, you need to use that info to DESCRIBE the discriminant. Be certain you know what "discriminant" signifies.

OpenStudy (mertsj):

\[(x+6)(4x+1)=4x^2+25x+6\]

OpenStudy (anonymous):

Ok, well I was given 4x^2-25X=-6., and I had to factor it. after this, it's asking me to use the discriminant to describe the type of solutions

OpenStudy (anonymous):

given ax^2 + bx + c if b^2 - 4ac > 0, there are two real solution if b^2 - 4ac < 0, there are no real solution if b^2 - 4ac = 0, there is only one solution

OpenStudy (mertsj):

The discriminant = \[b^2-4ac\]

OpenStudy (mathmale):

Oh. Please post the exact original question next time. As sourwing suggests, calculate the discriminant. Then decide which of the 3 possibilities for the discriminant value applies to your situation.

OpenStudy (anonymous):

3. Solve the equation below by factoring. 4x^2-25x=-6

OpenStudy (anonymous):

4) for each quadratic equation below, use the discriminant to describe the type of solutions